F(x) = 6x2 - 10x + 12


1
Expert's answer
2022-03-10T08:58:46-0500

Solution


Given that


"f\left( x \right) = 6{x^2} - 10x + 12\"


To solve, this "f\left( x \right) =0"


"6{x^2} - 10x + 12=0"


"3{x^2} - 5x + 6=0"


"x = \frac{{ - \left( { - 5} \right) \pm \sqrt {{{\left( { - 5} \right)}^2} - 4\left( 3 \right)\left( 6 \right)} }}{{2\left( 3 \right)}}\"


"x = \frac{{5 \pm \sqrt {25 - 72} }}{6}\"


"x = \frac{{5 \pm \sqrt { - 47} }}{6}\"


 "x = \frac{{5 + i\sqrt {47} }}{6}\" and "x = \frac{{5 - i\sqrt {47} }}{6}\"


Therefore, there are two imaginary solutions.


This has been shown by the following plot, which shows that the curve does not cross the x-axis. Hence no real solutions.